OpenStudy (anonymous):

Quadratic Formula: 0.05x^2-1x+1=0

OpenStudy (anonymous):

18.94427190999916 and 1.0557280900008425

OpenStudy (anonymous):

without calculator

OpenStudy (anonymous):

sry

OpenStudy (anonymous):

make your life easier and start with \[5x^2-100x+100=0\]

OpenStudy (anonymous):

then divide by \(5\) to get \[x^2-20x+20=0\]

OpenStudy (anonymous):

now either use the quadratic formula, or complete the square i will do the second method \[x^2-20x+20=0\] \[x^2-20x=-20\] \[(x-10)^2=-20+100=80\] \[x-10=\pm\sqrt{80}=\pm4\sqrt5\] \[x=10\pm4\sqrt5\]

OpenStudy (anonymous):

im trying to put it into quadratic formula and i got 10+-4√10

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok i think i know what the mistake would be

OpenStudy (anonymous):

you get \[a=1,b=-20,c=20\] and \[x=\frac{20\pm\sqrt{20^2-4\times 20}}{2}\] as a first step

OpenStudy (anonymous):

gives \[\frac{20\pm\sqrt{400-80}}{2}=\frac{20\pm\sqrt{320}}{2}\]

OpenStudy (anonymous):

now \[\sqrt{320}=8\sqrt{5}\] giving \[\frac{20\pm8\sqrt{5}}{2}=10\pm4\sqrt{5}\]

OpenStudy (anonymous):

okay that makes sense!!! Thank you soooo much!!!! :D